OpenStudy (anonymous):

A piece of fabric measures 40 by 38 inches. A triangular scarf with a height of 27 inches and a base of 30 inches is cut from the fabric. How much is left over? A. 355 in.^2 B. 557.5 in.^2 C. 1115 in.^2 D. 710 in.^2

OpenStudy (anonymous):

area of the fabric= 1520 sq inch area of triangle= .5*27*30 sq inches= 405 sq inches so, left over= 1520-405 sq inches=1115 sq inches (c)

Directrix (directrix):

Area of Cloth = L*W and Area of Triangular Flag = ½ b*h Area of Residual Fabric = 40*38 – ½(27)(30) assuming that the cloth is rectangular. Area = 1520 – 405 A = 1115 sq in